# Gravitational quasinormal modes of Dymnikova black holes

### v2

@@ I Introduction
QNMs of test fields in Dymnikova black holes have been investigated in four and higher dimensions in several recent works Konoplya and Zhidenko (2024a); Macêdo et al. (2024); Konoplya et al. (2023). However, to the best of our knowledge, no analogous analysis has yet been carried out for the most important case of gravitational perturbations within the framework of Asymptotically Safe gravity. In this paper we address this gap by computing the dominant QNMs of gravitational perturbations in the Dymnikova background. It is worth noting that grey-body factors and absorption cross-sections for gravitational perturbations of the Dymnikova solution were recently computed in Dubinsky (2025a). However, that analysis did not provide any results for the quasinormal frequencies.
@@ III Gravitational perturbations
Following this idea, Bouhmadi-López and collaborators Bouhmadi-López et al. (2020); Konoplya and Stashko (2025a) analyzed axial perturbations under the assumption that fluctuations aligned with the anisotropy do not contribute in the axial sector. This simplification is reminiscent of the treatment in Bronnikov et al. (2012), where the decoupling of scalar and axial gravitational modes was exploited in scalar–tensor models (see also Chen and Chen (2019)). Although such assumptions necessarily restrict the analysis, they have been widely adopted in related contexts and are generally expected to provide reliable results for small deviations from Schwarzschild geometry Konoplya and Stashko (2025b, a); Dubinsky (2025a); Bolokhov and Skvortsova (2025c). For Dymnikova black holes as solutions of the Einstein equations with the energy-momentum tensor corresponding to some density profile, the perturbation equations were considered in Dymnikova and Galaktionov (2005).
@@ IV Quasinormal modes
In the present analysis, our focus is on extracting the two lowest-lying quasinormal frequencies, $\omega_{0}$ and $\omega_{1}$. For this purpose, we employ the sixth or higher-order WKB formalism supplemented by Padé approximants with indices $\tilde{m}=\tilde{n}=3$ Konoplya et al. (2019). This refinement is known to provide optimal accuracy in many situations, as supported by a wide range of recent investigations Dubinsky (2025b); Skvortsova (2024b); Bronnikov and Konoplya (2020); Konoplya (2020); Bolokhov (2024a); Dubinsky (2025c); Skvortsova (2024c); Dubinsky and Zinhailo (2024); Malik (2024a); Churilova et al. (2021); Lütfüoğlu (2025b); Dubinsky (2025d); Malik (2025c); Dubinsky (2025e); Skvortsova (2024d); Bolokhov (2024b); Dubinsky and Zinhailo (2025). The convergence of results across these diverse studies highlights the reliability of the WKB–Padé scheme for accurately capturing the dominant part of the quasinormal spectrum.
@@
Notice that, by employing the correspondence between QNMs and grey-body factors Konoplya and Zhidenko (2024b); Lütfüoğlu (2025f); Malik (2025e); Bolokhov and Skvortsova (2025d); Han and Gwak (2025), the recently obtained grey-body factors in Dubinsky (2025a) can be reproduced from the numerical data presented in this work.
Taking into account the recent advances in gravitational-wave observations, which have achieved very high signal-to-noise ratios Abac et al. (2025), the estimated $4\%$ effect provides hope that such deviations may become observable. This raises the prospect of distinguishing classical black holes from their regular, quantum-corrected counterparts, even though the geometric modifications are confined primarily to a relatively small region near the event horizon.
@@
where $n$ is the overtone number. This correspondence holds for a broad class of black holes with single–peaked effective potentials, providing an intuitive geometrical interpretation of the quasinormal spectrum in terms of the motion of photons on unstable circular orbits. Nevertheless, there are a number of deviations from this relation, which are described in Khanna and Price (2017); Konoplya and Stuchlík (2017); Bolokhov (2024d); Konoplya (2023) and are usually related to gravitational or other non-minimally coupled perturbations. They may signal the breakdown of the eikonal approximation. Here, in Table VIII we study the high $\ell$ regime of QNMs for the near-extreme Dymnikova black hole. One can easily compute the proper characteristics of the null geodesics for the Dymnikov black hole and compare them with the numerical data presented in Table VIII. We conclude that the correspondence holds for gravitational perturbations in our case.
@@ V Conclusions
It is important to emphasize that the present study has been mainly restricted to the fundamental mode and the first overtone for which the WKB approach is reliable. For higher overtones, however, the situation is expected to be qualitatively different. Since overtones are strongly localized near the event horizon, they should be especially sensitive to the local geometric deformations induced by quantum corrections Konoplya and Zhidenko (2024c); Konoplya et al. (2022). To investigate this regime, one would need to resort to more precise techniques such as the Frobenius (Leaver) method Leaver (1985), which provides convergent series solutions and can capture the full spectrum beyond the eikonal approximation. At the same time, the metric, and consequently the master wave equation, does not have a rational form, so one has to transform or approximate the metric by some rational function for accurate calculations of higher overtones. Having in mind that the fundamental mode and possibly the first overtone are the only realistic candidates for detection in the near future experiments, we leave this analysis, as well as the inclusion of polar perturbations and possible extensions to higher dimensions, for future work.

### v3

@@ I Introduction
QNMs of test fields in Dymnikova black holes have been investigated in four and higher dimensions in several recent works Konoplya and Zhidenko (2024a); Macêdo et al. (2024); Konoplya et al. (2023). However, to the best of our knowledge, no analogous analysis has yet been carried out for the most important case of gravitational perturbations within the framework of Asymptotically Safe gravity. In this paper we address this gap by computing the dominant QNMs of gravitational perturbations in the Dymnikova background. It is worth noting that grey-body factors and absorption cross-sections for gravitational perturbations of the Dymnikova solution were recently computed in Dubinsky (2026). However, that analysis did not provide any results for the quasinormal frequencies.
@@ III Gravitational perturbations
Following this idea, Bouhmadi-López and collaborators Bouhmadi-López et al. (2020); Konoplya and Stashko (2025a) analyzed axial perturbations under the assumption that fluctuations aligned with the anisotropy do not contribute in the axial sector. This simplification is reminiscent of the treatment in Bronnikov et al. (2012), where the decoupling of scalar and axial gravitational modes was exploited in scalar–tensor models (see also Chen and Chen (2019)). Although such assumptions necessarily restrict the analysis, they have been widely adopted in related contexts and are generally expected to provide reliable results for small deviations from Schwarzschild geometry Konoplya and Stashko (2025b, a); Dubinsky (2026); Bolokhov and Skvortsova (2025c). For Dymnikova black holes as solutions of the Einstein equations with the energy-momentum tensor corresponding to some density profile, the perturbation equations were considered in Dymnikova and Galaktionov (2005).
@@ IV Quasinormal modes
In the present analysis, our focus is on extracting the two lowest-lying quasinormal frequencies, $\omega_{0}$ and $\omega_{1}$. For this purpose, we employ the sixth or higher-order WKB formalism supplemented by Padé approximants with indices $\tilde{m}=\tilde{n}=3$ Konoplya et al. (2019). This refinement is known to provide optimal accuracy in many situations, as supported by a wide range of recent investigations Dubinsky (2025a); Skvortsova (2024b); Bronnikov and Konoplya (2020); Konoplya (2020); Bolokhov (2024a); Dubinsky (2025b); Skvortsova (2024c); Dubinsky and Zinhailo (2024); Malik (2024a); Churilova et al. (2021); Lütfüoğlu (2025b); Dubinsky (2025c); Malik (2025c); Dubinsky (2025d); Skvortsova (2024d); Bolokhov (2024b); Dubinsky and Zinhailo (2025). The convergence of results across these diverse studies highlights the reliability of the WKB–Padé scheme for accurately capturing the dominant part of the quasinormal spectrum.
@@
More explicitly, in the light-cone coordinates $u=t-r_{*}$ and $v=t+r_{*}$ the wave equation takes the form
$\frac{\partial^{2}\Psi}{\partial u\,\partial v}+\frac{1}{4}V(r)\Psi=0.$ (20)
Discretizing this equation on a uniform $(u,v)$ grid with step size $\Delta$, one obtains the standard second-order accurate finite-difference scheme Gundlach et al. (1994)
$\Psi_{N}=\Psi_{W}+\Psi_{E}-\Psi_{S}-\frac{\Delta^{2}}{8}\,V\!\left(r_{*}\right)\left(\Psi_{W}+\Psi_{E}\right),$ (21)
where the subscripts denote the grid points $S=(u,v)$, $E=(u,v+\Delta)$, $W=(u+\Delta,v)$, and $N=(u+\Delta,v+\Delta)$. The effective potential $V(r)$ is evaluated at the center of the grid cell, corresponding to the tortoise coordinate $r_{*}(u+\Delta/2,v+\Delta/2)$. This scheme is conditionally stable and converges quadratically with respect to the grid spacing.
The initial data are specified by prescribing a compact Gaussian pulse on one of the null surfaces, for example,
$\Psi(u=u_{0},v)=\exp\!\left[-\frac{(v-v_{c})^{2}}{\sigma^{2}}\right],$ (22)
while setting $\Psi(v=v_{0},u)=0$ on the other surface. The subsequent evolution uniquely determines the waveform $\Psi(t,r_{*})$ at any fixed observation point outside the black hole. At intermediate times the signal is dominated by exponentially damped oscillations corresponding to the quasinormal ringing, while at late times it transitions to power-law tails.
To extract the dominant quasinormal frequencies, the time-domain signal in the ringdown regime is fitted to a superposition of damped exponentials,
$\Psi(t)\simeq\sum_{k=1}^{N}C_{k}e^{-i\omega_{k}t},$ (23)
using the Prony method. This technique reconstructs the complex frequencies $\omega_{k}$ by solving a linear prediction problem for a discrete time series sampled at equal time intervals. In practice, the fitting window is chosen such that the prompt response has decayed and the late-time tail has not yet become dominant. The stability of the extracted frequencies is verified by varying the fitting interval and the number of included modes. While the overtones are increasingly difficult to resolve in the time domain, the Prony method reliably captures the fundamental mode.
@@
Both the real oscillation frequency and the damping rate decrease as the quantum parameter $l_{\mathrm{cr}}$ increases, as can be seen for the fundamental mode and the first overtone across all multipole numbers $\ell=2$44 in Tables I–VII. At the same time, for a fixed $\ell$, the first overtone exhibits significantly stronger deviations from its Schwarzschild limit than the fundamental mode. For instance, in the $\ell=2$ case, the damping rate of the fundamental mode differs from the Schwarzschild value by approximately $4\%$, whereas the corresponding deviation for the first overtone exceeds $6\%$. The effect is even more pronounced for the real part of the frequency: while the shift of the fundamental mode is of the order of $1\%$, the first overtone shows a substantially larger deviation, reaching about $6\%$.
@@
Notice that, by employing the correspondence between QNMs and grey-body factors Konoplya and Zhidenko (2024b); Lütfüoğlu (2025f); Malik (2025e); Bolokhov and Skvortsova (2025d); Han and Gwak (2025), the recently obtained grey-body factors in Dubinsky (2026) can be reproduced from the numerical data presented in this work.
A deviation at the level of a few percent in the complex QNM frequency is, in principle, within the target range of black-hole spectroscopy, but its detectability depends primarily on the ringdown signal-to-noise ratio (SNR), the adopted ringdown model (single mode vs. multimode/overtones), and correlations with the remnant mass and spin. Current LIGO–Virgo–KAGRA (LVK) analyses typically constrain QNM frequency parameters at the ${\cal O}(10\%)$ level (or weaker) for individual events, with the achievable precision strongly varying from event to event and depending on the ringdown start time and modeling assumptions Abbott et al. (2025); Ghosh et al. (2021); Cotesta et al. (2022). Within this landscape, a $\sim 4\%$ shift in the damping rate and a $\sim 1\%$ shift in the real oscillation frequency should be regarded as challenging for present detectors in single-event measurements, but potentially accessible for particularly loud ringdowns and, more robustly, through the combination (stacking) of multiple detections, for which statistical uncertainties typically scale as $\propto 1/\sqrt{N}$ at fixed systematics.
Future third-generation (3G) ground-based detectors are expected to deliver substantially higher ringdown SNRs and enable percent-level spectroscopy for a sizable population of stellar-mass binary black-hole mergers Bhagwat et al. (2023); Berti et al. (2025). Forecasts indicate that, for favorable sources, multiple QNM parameters (frequency and damping time/quality factor) can be measured with a few-percent accuracy, making shifts at the ${\cal O}(1\%$$5\%)$ level observationally relevant in that era Bhagwat et al. (2023). Therefore, while the present results mainly quantify the size of the effect at the waveform level, they also provide a concrete target for future ringdown tests in high-SNR events and 3G-era population analyses.
Thus, taking into account the recent advances in gravitational-wave observations, which have achieved very high signal-to-noise ratios Abac et al. (2025), the estimated $4\%$ effect provides hope that such deviations may become observable. This raises the prospect of distinguishing classical black holes from their regular, quantum-corrected counterparts, even though the geometric modifications are confined primarily to a relatively small region near the event horizon.
@@
where $n$ is the overtone number. This correspondence holds for a broad class of black holes with single–peaked effective potentials, providing an intuitive geometrical interpretation of the quasinormal spectrum in terms of the motion of photons on unstable circular orbits. Nevertheless, there are a number of deviations from this relation, which are described in Khanna and Price (2017); Konoplya and Stuchlík (2017); Bolokhov (2024d) and are usually related to gravitational or other non-minimally coupled perturbations. Thus, in some theories with higher-curvature corrections—such as Gauss–Bonnet, Lovelock, or Einstein-Weyl gravity — the centrifugal term that dominates in the eikonal regime does not retain its standard form Konoplya and Zhidenko (2017); Konoplya et al. (2025a). Moreover, in asymptotically de Sitter spacetimes, an additional branch of QNMs emerges, corresponding to pure de Sitter modes modified by the presence of a black hole, for which the usual eikonal correspondence does not apply Konoplya et al. (2025b); Konoplya (2023). Such cases may signal the breakdown of the eikonal approximation. Here, in Table VIII we study the high $\ell$ regime of QNMs for the near-extreme Dymnikova black hole. One can easily compute the proper characteristics of the null geodesics for the Dymnikov black hole and compare them with the numerical data presented in Table VIII. We conclude that the correspondence holds for gravitational perturbations in our case.
This can also be shown analytically. An unstable null circular geodesic at a distance $r=r_{c}$ satisfies the condition
$V^{\prime}(r)\bigl|_{r_{c}}=0,$
that leads to the following relation,
$2f_{c}-r_{c}f^{\prime}_{c}=0.$ (25)
Then, the above relation is used to find the radius of the circular orbit $r_{c}$.
The angular velocity
$\Omega=d\varphi/dt=\dot{\varphi}/\dot{t}$
can be found using the circular motion condition $\dot{r}=0$:
$\Omega_{c}=\frac{\sqrt{f_{c}}}{r_{c}}.$ (26)
Finally, the Lyapunov exponent can be found as follows Cardoso et al. (2009),
$\lambda=\sqrt{\frac{f_{c}(2f_{c}-f^{\prime\prime}_{c}r_{c}^{2})}{2r_{c}^{2}}}.$ (27)
By expanding the location of the maximum of the effective potential in powers of $1/\ell$, which governs the decay of gravitational perturbations, one immediately finds that it coincides with the radius of the circular null orbit $r_{c}$. Consequently, applying the eikonal limit of the first-order WKB approximation to the quasinormal frequency directly reproduces Eq. (24).
@@ V Conclusions
It is important to emphasize that the present study has been mainly restricted to the fundamental mode and the first overtone for which the WKB approach is reliable. For higher overtones, however, the situation is expected to be qualitatively different. Since overtones are strongly localized near the event horizon, they should be especially sensitive to the local geometric deformations induced by quantum corrections Konoplya and Zhidenko (2024c); Konoplya et al. (2022). Unfortunately, the WKB approximation is reliable only when the multipole number $\ell$ is larger than, or at most comparable to, the overtone number $n$, and it generally fails in the regime $n\gg\ell$. In this case, the imaginary part of the frequency becomes large and the corresponding turning points are widely separated in the complex plane, which undermines the accuracy of the WKB expansion. Although the WKB method is computationally efficient and highly effective for the lowest modes, it does not guarantee convergence order by order; instead, only asymptotic convergence can be expected. At the same time, the phenomenon of overtone outbursts occurs precisely in the regime where the overtone number appears to exceed the multipole number, so that the relevant physics is governed by the near-horizon geometry rather than by the potential peak dominated by the centrifugal, $\ell$-dependent term. However, even with the WKB method within its expected accuracy range, we see that the first overtone deviates from its Schwarzschild limit much stronger than the fundamental mode, indicating that there should be an outburst of overtones found for test fields in Konoplya et al. (2023). To investigate this regime, one would need to resort to more precise techniques such as the Frobenius (Leaver) method Leaver (1985), which provides convergent series solutions and can capture the full spectrum beyond the eikonal approximation. At the same time, the metric, and consequently the master wave equation, does not have a rational form, so one has to transform or approximate the metric by some rational function for accurate calculations of higher overtones. Having in mind that the fundamental mode and possibly the first overtone are the only realistic candidates for detection in the near future experiments, we leave this analysis, as well as the inclusion of polar perturbations and possible extensions to higher dimensions, for future work.